EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Conditional Probability & Bayes' Theorem · Question 35 of 125

Question 35

EasyOriginal question

P(A)=0.7P(A) = 0.7 and P(B∣A)=0.4P(B \mid A) = 0.4. What is P(A∩B′)P(A \cap B')?

Options for question 35

Show the answer and explanation
CorrectNot quiteAnswer C

Correct answer: C

Explanation

  1. Given AA, P(B′∣A)=1−0.4=0.6P(B' \mid A) = 1 - 0.4 = 0.6.

  2. P(A∩B′)=P(A) P(B′∣A)=0.7×0.6=0.42P(A \cap B') = P(A)\,P(B' \mid A) = 0.7 \times 0.6 = 0.42.

Worth noting

0.28 is P(A∩B)=0.7×0.4P(A \cap B) = 0.7 \times 0.4, with BB where B′B' was asked for.

Concept
The multiplication theorem
Typical time
60 s
Your time
—
The rule this testsThe multiplication theorem
Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.

Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.